OpenStudy (anonymous):

What is the factored form of 5x^2+19x-4?

OpenStudy (anonymous):

5 x -4 = -20 now we have to split +19 in two parts [b1 and b2] in such a way that the summation of b1+b2 given 19 and their product gives -20 so 19 = 20-1 such that 20 x (-1) = -20 hence 5x^2 +20x - x -4 = 0 now take common ... hope you can do from here

OpenStudy (anonymous):

If you weren't able to do it from there the factored form would be (x + 4) (5x - 1). Btw good job explaining @digitalmonk :)

OpenStudy (anonymous):

you have me confused 5x-4 =-20 where did that come from

OpenStudy (anonymous):

(5x-1)(x+4) is much easier

OpenStudy (anonymous):

I explained you the process. 5x^2 +20x - x -4 = 0 5x(x+4) -1(x+4) = 0 i.e. (x+4)(5x-1) = 0 [taking x+4 as common] so either x+4 =0 which means x = -4 or (5x-1) = 0 which means x = - 1/5